OpenStudy (anonymous):

find the area of an equilateral triangle with sides 10 cm. ?

OpenStudy (cwrw238):

the altitude = sqrt(10^2 - 5^2) = sqrt75 area = 0.5 * base* altitude = 0.5* 10 * sqrt75 = 43.3 sq cms

OpenStudy (anonymous):

are of equilateral triangle = s^2*(root3/4) = 10^2(root3/4) = 100(root3/4) = 25root3